Saturday 03/23/13

10:30 A – 6:20P, 9:30P

Embryo collection

  • Flip all fly cages
  • dispose of previous fly plates with new plate recycling and freezing system.
  • Fix embryos 9:30 PM (~3-14 hrs, mostly at raised at 29C).


  • working out parameters for TCEP imaging

Current working set:

  • 128×128 field of view.
  • 310 mW 647.
  • 1750 mA of 752.
  • 2 mW max 405.
  • 60Hz frame rate for both. 45K frames of 750 for BXC. maybe 60K frames for Cy5 BXC.
  • 14.2 TIRF (any higher 750 fails. 647 would definetely prefer a slightly higher TIRF angle).
  • should have had longer bleach fro 647. (surprisingly little blinking for the amount of bleach time (~7000 frames at 60Hz).


  • signal to noise actually looks much better in 750. The slide may have aged a bit poorly, or this buffer is less good for cy5.

Experiment planning

  • embryo staining tomorrow, image on Wed.
  • Order primers for rest of LC library for chromatin colors
  • emPCR of rest of LC library on Thursday (starting at noon).
  • Fab7 start biological replicates in two weeks (Brian will make more probe after his department retreat).
This entry was posted in Research Planning and tagged , , . Bookmark the permalink.